OpenStudy (kpoplove2289):

A 30in board is cut into 3 pieces so that the second piece is 4 times as long as the first piece and the third piece is 5 times as long as the first piece. If x represents the length of the first piece, find the lengths of all three pieces.

OpenStudy (raffle_snaffle):

30in = stuff

OpenStudy (raffle_snaffle):

1st piece will just be x 2nd ? 3rd? What do you think?

OpenStudy (kpoplove2289):

y and z?

OpenStudy (raffle_snaffle):

we don't need y and z. The x represents a dimension or a length.

OpenStudy (kpoplove2289):

I'm sorry. So, then what do you mean with "What do you think?"

OpenStudy (vuriffy):

Hello.

OpenStudy (kpoplove2289):

Hello!

OpenStudy (vuriffy):

Look at the relation stated in the passage: 4 times and 5 times of the x value.

OpenStudy (vuriffy):

30 inches as a whole.

OpenStudy (kpoplove2289):

Understood so far.

OpenStudy (vuriffy):

I took a random value, for x, x = 3. 4(3) = 12. 5(3) = 15. 3 + 12 + 15 = 30. Amazingly, this is the correct value. I have always had a unique ability to guess the value quite quickly.

OpenStudy (vuriffy):

3 pieces, added all together equals the whole.

OpenStudy (vuriffy):

9x + 3 = 30.

OpenStudy (vuriffy):

9x = 27 x = 3. For Algebra stance.

OpenStudy (vuriffy):

Then just plug in 3 for 4(x) and 5(x) and see if it adds up to 30, which it does.

OpenStudy (raffle_snaffle):

x = inches

OpenStudy (vuriffy):

I apologize, raffle. I had to share when I found the answer on my first guess, haha.

OpenStudy (raffle_snaffle):

4 times, means 4 times greater than x amount.

OpenStudy (raffle_snaffle):

No worries, more help the better as long as OP is learning I am fine.

OpenStudy (vuriffy):

Kpop, understand everything we said? Need any further explanation?

OpenStudy (kpoplove2289):

Wait, did you randomly just choose 3? Because, what I'm more confused about is trying to make it into an equation then solving for x.

OpenStudy (vuriffy):

Yes, I randomly choice 3, myself, but when I looked at the problem further I see that there is 4x and 5x and 3 pieces. So I constructed a equation like this: 9x + 3 = 30 (the total inches)

OpenStudy (vuriffy):

Then solved for x.

OpenStudy (vuriffy):

There's always more than one way to solve a puzzle.

OpenStudy (kpoplove2289):

Would you know how to put it into an equation without guessing numbers? Is there a formula I can use?

OpenStudy (vuriffy):

I didn't use a formula, more so, I looked at what the passage gave me. The equation I made wasn't a guess. 4x + 5x + 3 = 30. All from the passage.

OpenStudy (kpoplove2289):

i meant without guessing a number

OpenStudy (vuriffy):

9x + 3 = 30 9x = 27 x = 3 3 + 4(3) + 5(3) 3 + 12 + 15 = 30 30 = 30.

OpenStudy (vuriffy):

I never guessed a number in what I just stated.

OpenStudy (vuriffy):

All this information came from what you stated.

OpenStudy (kpoplove2289):

Oh!! Because of the three pieces, right?

OpenStudy (vuriffy):

Yes.

OpenStudy (vuriffy):

I only combined like-terms with the 4x and 5x, also.

OpenStudy (kpoplove2289):

So, for example if it was cut into 5 pieces, I can use the number 5 just like what you did with 3?

OpenStudy (vuriffy):

If need be, then I would hope so, because that's how I saw it, and it worked out just fine. So, the chances of that happening has a good value, but it doesn't mean it will work every time.

OpenStudy (vuriffy):

Let's see, I will make an example:

OpenStudy (vuriffy):

75 = 5 + 14x (x) + 2x + 3x + 4x + 5x

OpenStudy (vuriffy):

For 5 pieces.

OpenStudy (vuriffy):

70 = 14x x = 5

OpenStudy (vuriffy):

So, it does work in that case also. So, it seems to work in more than one case.

OpenStudy (vuriffy):

5 + 10 + 15 + 20 + 25 = 75, by the way.

OpenStudy (vuriffy):

1(5) + 2(5) + 3(5) + 4(5) + 5(5) = 75.

OpenStudy (kpoplove2289):

Oh, okay. I think this is much easier than how I tried learning it in class. It was more detailed in class.

OpenStudy (vuriffy):

Always remember there is more than one way to solve a puzzle, no matter the difficulty. (:

OpenStudy (kpoplove2289):

Haha, I will. Thanks so much. I really appreciate you helping me out.

OpenStudy (vuriffy):

Anytime, medal if you wish. You have a goodnight/good day! (:

OpenStudy (kpoplove2289):

Of course. Thank you. You too :)